OpenStudy (anonymous):

Solve the quadratic function x+5/3x+2-4-2x/3x^2+8x+4=x+4/x+2

OpenStudy (anonymous):

Okay so i already i already know to factor the 3x^2 part, and then find the LCD, but my end answer isn't corresponding with the book and i have no idea how to do it. HELP

OpenStudy (campbell_st):

do some factorising \[\frac{x+5}{3x+2} - \frac{2(2-x)}{(x+2)(3x+2)} = \frac{x+4}{x+2}\] the common denominator is (x+2)(3x+2) the getting everything over a common denominator \[\frac{(x+5)(x+2)-(x-2)}{(3x+2)(x+2)} =\frac{(x+4)(3x+2)}{(3x+2)(x+2)}\] I'd expand and simplify the numerators \[\frac{x^2 + 7x + 10 -4+2x}{(3x+2)(x+2)} = \frac{(x+4)(x+2)}{(3x+2)(x+2)}\]